• 10. A rocket ship has a mass of 1.50 x 105 kg. Its relativistic momentum is 3.00 x 1012 kg•m/s. How fast is the rocket ship traveling?


  • How fast would a meterstick be moving so that its length will be observed to shrink to half its original length?


  • How much work must be done to a proton to accelerate it from rest to 0.998c.


  • A person on Earth observes a space ship approaching from the right with a velocity of 0.75c and a rocket ship approaching from the left at 0.50c. What is the relative velocity of the two ships as measured by a passenger in each of them?
1
Expert's answer
2022-05-11T13:53:23-0400

"u=\frac{v_1+v_2}{1+\frac{v_1v_2}{c^2}}=0.91c,"

"0.5l=l\sqrt{1-\frac{v^2}{c^2}},\implies"

"v=\frac {\sqrt 3} 2c=0.87c."